# Near–Real-Time Conflict-Related Fire Detection Using Unsupervised Deep Learning and Satellite Imagery

### v2

@@ 4.2 Preprocessing and Spectral Alignment
To mitigate sensor-specific radiometric bias and atmospheric variability, we implement a dual-stage normalization pipeline. Initially, to align the 3-meter imagery with the 10-meter reference, we perform Spectral Normalization via Major Axis (MA) Linear Regression Chastain et al. (2019). For each overlapping spectral band $b$, the calibrated 3-meter reflectance $x_{\text{cal}}$ is derived as:
@@ 4.3 Evaluation and Metrics
Training: The model was trained using the WorldFloods dataset (Mateo-Garcia et al., 2021) derived from Sentinel-2 multi-spectral imagery (Fletcher, 2012). The dataset includes four classes of disaster-related surface changes, including fire burn scars. The dataset was selected for its high variance in spectral and textural patterns. Training on diverse disaster-related surface changes encourages the VAE to learn generalized, robust visual primitives rather than overfitting to specific biome characteristics. This capability enables the model to effectively reconstruct nominal arid backgrounds in Sudan and highlight fire-related deviations as anomalies, despite the ecological domain shift. Training was performed on 233 scenes, with a held-out validation set of 19 scenes used to assess convergence and prevent overfitting. Training and validation were conducted on a machine equipped with one NVIDIA H100 GPU, 112 Intel CPU cores, and 2 TB of system memory.
@@
Inference: For conflict monitoring in Sudan, radiometrically calibrated 4-band ortho-analytic imagery with red, green, blue, and near-infrared bands was partitioned into $32\times 32$ pixel tiles. Corresponding pre- and post-incident tiles from the same spatial locations were passed through the encoder to obtain latent embeddings. Anomaly detection is formulated as change detection in latent space rather than reconstruction-based scoring. Anomaly scores were computed using the cosine distance between the latent representations of each tile pair, capturing deviations from nominal surface conditions learned during training. This representation-based comparison reduces sensitivity to absolute radiometric differences and sensor noise, enabling robust detection of fire-affected areas without requiring labeled training data.
@@
Baseline comparison: We implemented a tile-wise cosine distance baseline for comparison. Tiles of size $32\times 32$ pixels were flattened into vectors of raw pixel values across all four bands. Cosine distances between corresponding pre- and post-incident tiles were computed to yield continuous anomaly scores. While spectral indices such as NDVI or BAI could utilize the available visible-NIR bands, they impose fixed arithmetic assumptions about surface reflectance that vary by biome. We selected the baseline to isolate the benefit of learned latent representations: by holding preprocessing, tile geometry, temporal pairing, and the similarity metric constant, any performance difference reflects the contribution of the VAE encoder, independent of heuristic spectral thresholds, allowing evaluation of whether change detection benefits from representation learning beyond direct pixel-space comparison. Baseline scores were thresholded at the 95th percentile of distances across nominal scenes and aggregated into scene-level prediction maps, enabling direct comparison with VAE-based anomaly scores derived from latent embeddings.
@@
Fire incident labeling: Ground-truth labels were curated as part of the SCO initiative, which analyzes information collected from human rights organizations, ACLED, and field interviews (Etefa, 2026; Rothbart et al., 2024a). Tiles were labeled fire-affected if post-incident imagery revealed burn scars or active fire. Each scene was reviewed by at least three analysts, with discrepancies resolved by consensus, yielding inter-annotator agreement above 90%. To mitigate circularity and spatial uncertainly, fire incidents were further validated using FIRMS and VIIRS datasets and cross-referenced with independent reports (Howarth et al., 2024a, b). Labels were used solely for retrospective evaluation and not for training, threshold optimization, or model selection. Visual interpretation of PlanetScope imagery was restricted specifically to the spatial delineation of burn scars for these pre-verified incidents. This ensures that the model is evaluated on its ability to segment confirmed damage, rather than on incidents defined solely by the imagery itself.
@@
Metrics: Model performance was quantified using the area under the precision–recall curve (AUPRC), which provides a threshold-independent measure of how effectively anomaly scores rank fire-affected tiles above non-fire tiles. AUPRC is preferred over the ROC curve in this setting because fire-affected tiles constitute a small fraction of each scene, and precision–recall analysis provides a more informative evaluation under severe class imbalance. Continuous anomaly scores, derived from cosine distances computed either in the VAE latent embedding space or directly in pixel space for the baseline method, were evaluated against the curated fire labels. Precision and recall were computed by sweeping thresholds across all observed scores to generate binary predictions, producing a precision–recall curve summarizing the trade-off between true and false positives. In addition, representative thresholds were used to report precision, recall, and F1-score for interpretable point estimates. Beyond quantitative metrics, model outputs were also assessed qualitatively by visualizing the spatial distribution of anomalous tiles, enabling inspection of spatial coherence and localization and illustrating correspondence with observed fire patterns in the imagery.
@@
Near–real-time detection: We define “near–real-time” detection as producing fire anomaly outputs within 24–30 hours of image acquisition. This timeframe includes data acquisition latency (approximately 18–20 hours for PlanetScope imagery), preprocessing and tiling (spectral standardization, log transforms, and $32\times 32$ tile generation in 1–2 hours), and model inference plus postprocessing (reconstruction, scoring, and visualization in 1–2 hours per scene on GPU-enabled hardware). Together, these stages enable rapid situational awareness for monitoring fire-affected areas in conflict settings.
@@ 4.4 Imagery details
Planet operates Earth-imaging constellations: PlanetScope (PS), RapidEye (RE), and SkySat (SS). Imagery formats serve various use cases, including deep learning, disaster response, precision agriculture, and temporal image analytics. Planet offers three product lines for PlanetScope imagery: Basic Scene, Ortho Scene, and Ortho Tile. The Basic Scene product is a scaled Top of Atmosphere Radiance (at sensor) and sensor-corrected product, suitable for advanced image processing and geometric correction. Ortho Tiles are multiple orthorectified scenes in a single strip, divided according to a defined grid. Ortho Scenes represent single-frame image captures with post-processing, removing terrain distortions and suitable for cartographic purposes. Ortho Scenes are delivered as visual (RGB) and analytic products (Marta, 2018). In this work, we utilized radiometrically calibrated 4-band ortho-analytic imagery obtained from the PS2 instrument, which captures red, green, blue, and near-infrared channels.
Table 1: Details of satellite imagery used for the study area.
Place Area Incident Before-imagery After-imagery
($km^{2}$) date date date
Gandahar Market 3.072 May 25, 2024 May 13, 2024 May 25, 2024
El Fasher 4.032 May 06-07, 2024 May 02, 2024 May 11, 2024
Muqrin 3.168 Apr 13, 2024 Mar 01, 2024 Apr 13, 2024
Jaranga 1.536 Apr 13, 2024 Mar 01, 2024 Apr 13, 2024
Sarafaya 1.44 Apr 13, 2024 Mar 01, 2024 Apr 13, 2024
@@ 4.5 Incident data
The SCO methodology for collecting incident data uses a multi-phase approach combining secondary research and original data collection. Researchers first examine reports from organizations including Human Rights Watch, the World Health Organization, and the United Nations. Additionally, Armed Conflict Location & Event Data (ACLED) catalog is utilized as another complementary source to consolidate the incidents dataset. These sources document civilian targeting and infrastructure destruction, which the SCO then aims to enhance with further analysis or more current accounts from the impacted areas in Sudan. Social media provides direct witness reports and real-time updates for incident verification. By combining these varied streams, the SCO researchers confirm the validity of specific incidents. Furthermore, Sudan Human Rights Hub (SHRH) conducted interviews of individuals on the ground affected by the attacks. Testimony was given only after the research participants received full information about the purpose, scope and use of their information, followed by their consent. Detailed methodology and technical specifications of the incident collection methodology are reported in (Masri et al., 2025)
ACLED catalog compiles and curates a dataset of individual political disorder events across the world. Its focus is on political violence and protest, and it monitors the activity of a variety of conflict actors and armed groups, as well as multiple forms of unrest. Unlike other datasets that use machine-led coding, ACLED relies on a researcher-led data collection process that allows for the incorporation of conflict and protest events from a wide range of sources, including local-language media, conflict observatories and other situation reports, which mitigates over-reliance on international news agencies (Carboni and Raleigh, 2024)
Since the scope of the SCO reporting is broader, ranging from infrastructure damages to civil rights violations to gender violence, we manually selected the incidents identified as destruction attributed to fires in the conflict. The rationale for selecting such incidents was that satellite imagery can capture and validate the visual signatures of fire-related destruction reported by the SCO.
@@ 5 Results
We evaluate the effectiveness of the proposed deep learning–based fire detection approach by comparing model outputs against a baseline that computes cosine distance between pre- and post-incident image tiles after identical preprocessing. Performance was assessed across five case studies in Sudan using qualitative visual inspection of prediction maps and quantitative metrics. To quantify uncertainty and robustness, we computed 95% confidence intervals for all metrics via 1,000 bootstrap resamples of tiles within each scene. Additionally, paired Wilcoxon signed-rank tests were applied to AUPRC values obtained from identical bootstrap resamples of tiles for the baseline and proposed method, ensuring matched, nonparametric comparison under identical sampling conditions. In addition to absolute performance metrics, we report Cohen’s $d$ to quantify the magnitude of improvement of the proposed VAE-based approach relative to the baseline. Figures 3 and 4 present visual comparisons for Jaranga and El Fasher, respectively.
@@
Figure 3: Comparison of baseline with prediction in Jaranga.
@@
Figure 4: Comparison of baseline with prediction in El Fasher.
Results indicate that the baseline method tends to replicate fine-grained details from the input imagery and is highly sensitive to noise, misregistration, and small radiometric or geometric variations. In contrast, the proposed approach captures higher-level latent features, facilitating more robust delineation of fire-affected areas from background changes.
@@
Figures 5, 6, 7, 8 and 9 provide qualitative results across all study areas. The model successfully detects active fires and associated smoke plumes in Gandahar Market in Figure 5. Similarly, Figure 6 demonstrates the model’s ability to identify burn scars following fire incidents. We note that large areas of El Fasher are burned after multiple successive fire incidents, leaving extensive swaths of charred surfaces. Even in relatively complex images, the model performs well in delineating severely burned areas. Figure 7 shows that multiple simultaneous fires are detected in Jaranga, while previously impacted locations are correctly identified in Muqrin (Figure 8). Finally, Figure 9 illustrates accurate delineation of recently burned surfaces in Sarafaya. Across all cases, predicted fire-affected regions exhibit strong spatial coherence and align well with independently verified ground-truth labels.
@@ 5.1 Quantitative Performance
Tables 26 summarize bootstrapped metrics for each location and imagery configuration. Across all sites, the proposed VAE-based approach consistently outperformed the baseline. Median relative AUPRC improvements ranged from 18% to 40%, corresponding to moderate-to-large effect sizes (Cohen’s $d$ = 0.8–1.5). Relative improvement is computed using median values as $(\mathrm{AUPRC}_{\text{pred}}-\mathrm{AUPRC}_{\text{base}})/\mathrm{AUPRC}_{\text{base}}$. Paired statistical tests confirmed that differences were statistically significant ($p<0.01$) for all locations. Improvements are particularly pronounced in complex urban and peri-urban settings such as El Fasher and Jaranga, where fires are spatially fragmented and embedded in heterogeneous backgrounds. In Muqrin and Sarafaya, which exhibit clearer fire signatures, the model achieves AUPRC values exceeding 0.9. We note that the confidence interval widths vary modestly across sites due to differences in tile counts and spatial heterogeneity, but remain narrow owing to the large number of tiles per scene.
@@
Table 2: Gandahar Market: Bootstrapped metrics (median and 95% CI) and Cohen’s $d$.
Imagery Metric Baseline Prediction Cohen’s $d$
4-band AUPRC 0.584 (0.562–0.605) 0.747 (0.729–0.765) 1.12
Precision 0.63 (0.61–0.65) 0.71 (0.69–0.73) 0.87
Recall 0.55 (0.53–0.57) 0.74 (0.72–0.76) 1.33
F1 0.59 (0.57–0.61) 0.72 (0.70–0.74) 1.06
8-band AUPRC 0.619 (0.597–0.640) 0.825 (0.807–0.841) 1.38
Precision 0.66 (0.64–0.68) 0.76 (0.74–0.78) 0.96
Recall 0.57 (0.55–0.59) 0.77 (0.75–0.79) 1.40
F1 0.61 (0.59–0.63) 0.76 (0.74–0.78) 1.20
Time AUPRC 0.599 (0.578–0.618) 0.794 (0.775–0.811) 1.24
series Precision 0.64 (0.62–0.66) 0.73 (0.71–0.75) 0.89
Recall 0.56 (0.54–0.58) 0.75 (0.73–0.77) 1.35
F1 0.59 (0.57–0.61) 0.74 (0.72–0.76) 1.13
Table 3: El Fasher: Bootstrapped metrics (median and 95% CI) and Cohen’s $d$.
Imagery Metric Baseline Prediction Cohen’s $d$
4-band AUPRC 0.447 (0.430–0.464) 0.695 (0.678–0.710) 1.45
Precision 0.52 (0.50–0.54) 0.63 (0.61–0.65) 0.88
Recall 0.48 (0.46–0.50) 0.70 (0.68–0.72) 1.35
F1 0.50 (0.48–0.52) 0.66 (0.64–0.68) 1.15
8-band AUPRC 0.346 (0.331–0.360) 0.471 (0.455–0.486) 0.90
Precision 0.44 (0.42–0.46) 0.53 (0.51–0.55) 0.72
Recall 0.38 (0.36–0.40) 0.48 (0.46–0.50) 0.81
F1 0.41 (0.39–0.43) 0.50 (0.48–0.52) 0.78
Time AUPRC 0.407 (0.392–0.422) 0.646 (0.630–0.662) 1.32
series Precision 0.49 (0.47–0.51) 0.61 (0.59–0.63) 0.86
Recall 0.43 (0.41–0.45) 0.65 (0.63–0.67) 1.22
F1 0.45 (0.43–0.47) 0.63 (0.61–0.65) 1.05
Table 4: Muqrin: Bootstrapped metrics (median and 95% CI) and Cohen’s $d$.
Imagery Metric Baseline Prediction Cohen’s $d$
4-band AUPRC 0.673 (0.655–0.690) 0.929 (0.915–0.941) 1.54
Precision 0.71 (0.69–0.73) 0.90 (0.88–0.92) 1.12
Recall 0.63 (0.61–0.65) 0.94 (0.92–0.95) 1.68
F1 0.67 (0.65–0.69) 0.92 (0.90–0.94) 1.44
8-band AUPRC 0.691 (0.673–0.708) 0.935 (0.921–0.946) 1.48
Precision 0.73 (0.71–0.75) 0.91 (0.89–0.93) 1.09
Recall 0.65 (0.63–0.67) 0.94 (0.92–0.95) 1.61
F1 0.68 (0.66–0.70) 0.93 (0.91–0.94) 1.41
Time AUPRC 0.683 (0.665–0.700) 0.938 (0.924–0.950) 1.52
series Precision 0.72 (0.70–0.74) 0.92 (0.90–0.94) 1.16
Recall 0.64 (0.62–0.66) 0.95 (0.93–0.96) 1.64
F1 0.68 (0.66–0.70) 0.93 (0.91–0.95) 1.44
Table 5: Jaranga: Bootstrapped metrics (median and 95% CI) and Cohen’s $d$.
Imagery Metric Baseline Prediction Cohen’s $d$
4-band AUPRC 0.417 (0.401–0.433) 0.630 (0.612–0.646) 1.21
Precision 0.50 (0.48–0.52) 0.61 (0.59–0.63) 0.87
Recall 0.43 (0.41–0.45) 0.66 (0.64–0.68) 1.33
F1 0.46 (0.44–0.48) 0.63 (0.61–0.65) 1.12
8-band AUPRC 0.513 (0.495–0.531) 0.722 (0.703–0.738) 1.15
Precision 0.56 (0.54–0.58) 0.68 (0.66–0.70) 0.91
Recall 0.49 (0.47–0.51) 0.70 (0.68–0.72) 1.22
F1 0.52 (0.50–0.54) 0.69 (0.67–0.71) 1.09
Time AUPRC 0.474 (0.457–0.491) 0.671 (0.652–0.688) 1.18
series Precision 0.53 (0.51–0.55) 0.65 (0.63–0.67) 0.89
Recall 0.46 (0.44–0.48) 0.68 (0.66–0.70) 1.26
F1 0.49 (0.47–0.51) 0.66 (0.64–0.68) 1.12
Table 6: Sarafaya: Bootstrapped metrics (median and 95% CI) and Cohen’s $d$.
Imagery Metric Baseline Prediction Cohen’s $d$
4-band AUPRC 0.686 (0.668–0.704) 0.918 (0.905–0.932) 1.50
Precision 0.73 (0.71–0.75) 0.91 (0.89–0.93) 1.10
Recall 0.66 (0.64–0.68) 0.94 (0.92–0.95) 1.62
F1 0.69 (0.67–0.71) 0.92 (0.90–0.94) 1.42
8-band AUPRC 0.676 (0.658–0.693) 0.926 (0.913–0.939) 1.55
Precision 0.72 (0.70–0.74) 0.92 (0.90–0.94) 1.19
Recall 0.67 (0.65–0.69) 0.95 (0.93–0.96) 1.64
F1 0.69 (0.67–0.71) 0.93 (0.91–0.95) 1.45
Time AUPRC 0.694 (0.676–0.711) 0.929 (0.915–0.942) 1.52
series Precision 0.74 (0.72–0.76) 0.92 (0.90–0.94) 1.12
Recall 0.67 (0.65–0.69) 0.95 (0.93–0.96) 1.61
F1 0.70 (0.68–0.72) 0.93 (0.91–0.95) 1.42

### v3

@@ 4.2 Preprocessing and Spectral Alignment
To mitigate sensor-specific radiometric bias and atmospheric variability, we implement a dual-stage normalization pipeline. Initially, to align the 3-meter imagery with the 10-meter reference, we perform Spectral Normalization via Major Axis (MA) Linear Regression [16]. For each overlapping spectral band $b$, the calibrated 3-meter reflectance $x_{\text{cal}}$ is derived as:
@@ 4.3 Evaluation and Metrics
Training: The model was trained using the WorldFloods dataset [45] derived from Sentinel-2 multi-spectral imagery [26]. The dataset includes four classes of disaster-related surface changes, including fire burn scars. The dataset was selected for its high variance in spectral and textural patterns. Training on diverse disaster-related surface changes encourages the VAE to learn generalized spatial and structural features rather than overfitting to specific biome characteristics. This capability enables the model to effectively reconstruct nominal arid backgrounds in Sudan and highlight fire-related deviations as anomalies, despite the ecological domain shift. Training was performed on 233 scenes, with a held-out validation set of 19 scenes used to assess convergence and prevent overfitting. Training and validation were conducted on a machine equipped with one NVIDIA H100 GPU, 112 Intel CPU cores, and 2 TB of system memory.
@@
Inference: For conflict monitoring in Sudan, radiometrically calibrated 4-band ortho-analytic imagery with red, green, blue, and near-infrared bands was partitioned into $32\times 32$ pixel tiles. Corresponding pre- and post-incident tiles from the same spatial locations were passed through the encoder to obtain latent embeddings. Anomaly detection is formulated as change detection in latent space rather than reconstruction-based scoring. Anomaly scores were computed using the cosine distance between the latent representations of each tile pair, capturing deviations from nominal surface conditions learned during training. This representation-based comparison reduces sensitivity to absolute radiometric differences and sensor noise, facilitating the sensitive detection of fire-affected areas without requiring labeled training data.
@@
Comparison with conventional methods: We implemented a comparative evaluation using standard pixel-level change detection approaches, including cosine distance, canonical variates analysis (CVA) [41], and Iteratively Reweighted Multivariate Alteration Detection (IR-MAD) [50]. Cosine distances were computed between corresponding pre- and post-incident tiles to yield continuous anomaly scores, while CVA and IR-MAD provided change scores in original spectral space and statistically derived change significance measures, respectively. While spectral indices such as NDVI or BAI could leverage visible–NIR bands, they rely on fixed arithmetic assumptions about surface reflectance that vary across land cover types. By holding preprocessing, tile geometry, and temporal pairing constant, differences in performance isolate the contribution of the VAE-based latent representations, independent of heuristic thresholds. Scores from comparison methods were thresholded at the 95th percentile of dissimilarity values across nominal scenes and aggregated into scene-level prediction maps, enabling direct comparison with VAE-based anomaly scores derived from latent embeddings.
@@
Fire incident labeling: Ground-truth labels were curated as part of the SCO initiative, which analyzes information collected from human rights organizations, ACLED, and field interviews [25, 53]. Tiles were labeled fire-affected if post-incident imagery revealed burn scars or active fire. Each scene was reviewed by at least three analysts, with discrepancies resolved by consensus, yielding inter-annotator agreement above 90%. To mitigate circularity and spatial uncertainly, fire incidents were further validated using FIRMS and VIIRS datasets and cross-referenced with independent reports [32, 33]. Labels were used solely for retrospective evaluation and not for training, threshold optimization, or model selection. Visual interpretation of PlanetScope imagery was limited to the spatial delineation of burn scars for independently reported and verified incidents. This approach ensures that the model is evaluated on its ability to segment confirmed damage, rather than on incidents identified directly from satellite imagery.
@@
Metrics: Model performance was quantified using the area under the precision–recall curve (AUPRC), which provides a threshold-independent measure of how effectively anomaly scores rank fire-affected tiles above non-fire tiles. AUPRC is preferred over the ROC curve in this setting because fire-affected tiles constitute a small fraction of each scene, making precision–recall curves more informative under severe class imbalance. Continuous anomaly scores were derived from cosine distances in the VAE latent space and from pixel-space distance or transformation-based measures for the comparison methods, and were evaluated against curated fire labels. Precision and recall were computed by sweeping thresholds across all observed scores to generate binary predictions, producing a precision–recall curve summarizing the trade-off between true and false positives. Representative thresholds were selected to compute precision, recall, and F1-score, reported as median point estimates across folds. Beyond quantitative metrics, model outputs were assessed qualitatively by visualizing the spatial distribution of anomalous tiles, allowing inspection of spatial coherence and alignment with observed fire patterns.
@@
Near–real-time detection: We define “near–real-time” detection as producing fire anomaly outputs within 24–30 hours of image acquisition. This timeframe assumes ideal optical conditions, with cloud-free and accurately co-registered imagery, and includes data acquisition latency (approximately 18–20 hours for PlanetScope imagery), preprocessing and tiling (spectral standardization, log transforms, and $32\times 32$ tile generation in 1–2 hours), and model inference plus postprocessing (reconstruction, scoring, and visualization in 1–2 hours per scene on GPU-enabled hardware). While these stages enable rapid situational awareness under favorable conditions, actual operational latency may be longer in conflict zones due to cloud cover, image misregistration, or tasking constraints, which were not explicitly stress-tested in this study. Therefore, this definition serves primarily as a benchmark for method evaluation, and real-world deployment may require additional buffering to account for image availability and quality limitations.
@@ 4.4 Imagery details
Planet operates Earth-imaging constellations: PlanetScope (PS), RapidEye (RE), and SkySat (SS). Imagery formats serve various use cases, including deep learning, disaster response, precision agriculture, and temporal image analytics. Planet offers three product lines for PlanetScope imagery: Basic Scene, Ortho Scene, and Ortho Tile. The Basic Scene product is a scaled Top of Atmosphere Radiance (at sensor) and sensor-corrected product, suitable for advanced image processing and geometric correction. Ortho Tiles are multiple orthorectified scenes in a single strip, divided according to a defined grid. Ortho Scenes represent single-frame image captures with post-processing, removing terrain distortions and suitable for cartographic purposes. Ortho Scenes are delivered as visual (RGB) and analytic products [42]. In this work, we utilized radiometrically calibrated 4-band ortho-analytic imagery obtained from the PS2 instrument, which captures red, green, blue, and near-infrared channels. For time-series analyses, three pre-incident images were used to capture temporal changes, and the 8-band imagery included Coastal Blue, Blue, Green I, Green, Yellow, Red, Red Edge, and Near-Infrared bands.
@@ 4.5 Incident data
The SCO methodology for collecting incident data uses a multi-phase approach combining secondary research and original data collection. Researchers first examine reports from organizations including Human Rights Watch, the World Health Organization, and the United Nations. Additionally, Armed Conflict Location & Event Data (ACLED) catalog is utilized as another complementary source to consolidate the incidents dataset. These sources document civilian targeting and infrastructure destruction, which the SCO then aims to enhance with further analysis or more current accounts from the impacted areas in Sudan. Social media provides direct witness reports and real-time updates for incident verification. By combining these varied streams, the SCO researchers confirm the validity of specific incidents. Furthermore, Sudan Human Rights Hub (SHRH) conducted interviews of individuals on the ground affected by the attacks. Testimony was given only after the research participants received full information about the purpose, scope and use of their information, followed by their consent. Detailed methodology and technical specifications of the incident collection methodology are reported in [44].
ACLED catalog compiles and curates a dataset of individual political disorder events across the world. Its focus is on political violence and protest, and it monitors the activity of a variety of conflict actors and armed groups, as well as multiple forms of unrest. Unlike other datasets that use machine-led coding, ACLED relies on a researcher-led data collection process that allows for the incorporation of conflict and protest events from a wide range of sources, including local-language media, conflict observatories and other situation reports, which mitigates over-reliance on international news agencies [15].
SCO reports cover a wide range of events, including infrastructure damage, civil rights violations, and gender-based violence. For this study, we focused on incidents attributed to fire, as these produce visual signatures that can be captured and validated in satellite imagery. To ensure suitable data for evaluation, incidents were further filtered to include only those with cloud-free, accurately co-registered PlanetScope imagery within 24 hours of the reported event. From this subset, five case studies were selected for near–real-time benchmarking. This workflow provides case studies that are representative of fire-affected areas while maintaining consistent data quality for quantitative assessment.
@@ 5 Results
We evaluated the effectiveness of the proposed deep learning–based fire detection approach by comparing model outputs against CVA, cosine distance, and IR-MAD between pre- and post-incident images after identical preprocessing. Performance was assessed across five case studies in Sudan using qualitative visual inspection of prediction maps and quantitative metrics. To quantify uncertainty and robustness, we estimated 95% confidence intervals (CIs) for all performance metrics using 1,000 bootstrap resamples of tiles within each scene. Additionally, paired Wilcoxon signed-rank tests were applied to AUPRC values obtained from identical bootstrap resamples of tiles across methods, ensuring matched, nonparametric comparison under identical sampling conditions.
Across the five case studies, the VAE-based approach exhibits varying levels of sensitivity and robustness depending on scene characteristics. In scenarios with strong burn signals, the model achieves high recall and clearly delineates fire-affected areas, while in high-noise or degraded conditions, spectral limitations can reduce accuracy.
@@
Figure 3: Comparison of cosine distance with prediction in Jaranga.
@@
Figure 4: Comparison of cosine distance with prediction in El Fasher.
@@
Figures 3 and 4 present visual comparisons with cosine distance for Jaranga and El Fasher, respectively. Results indicate that cosine distance preserves fine-grained details from the input imagery but is sensitive to noise, misregistration, and minor radiometric or geometric variations. In contrast, the proposed VAE-based approach captures higher-level latent features, facilitating the delineation of fire-affected areas relative to background changes.
@@
Figures 5, 6, 7, 8 and 9 provide qualitative results across all study areas. The model successfully detects active fires and associated smoke plumes in Gandahar Market in Figure 5. Similarly, Figure 6 demonstrates the model’s ability to identify burn scars following fire incidents. We note that large areas of El Fasher are burned after multiple successive fire incidents, leaving extensive swaths of charred surfaces. Even in relatively complex images, the model is generally able to delineate severely burned areas with reasonable accuracy. Figure 7 shows that multiple simultaneous fires are detected in Jaranga, while previously impacted locations are correctly identified in Muqrin (Figure 8). Finally, Figure 9 indicates accurate delineation of recently burned surfaces in Sarafaya. Across cases, predicted fire-affected regions exhibit spatial coherence and demonstrate good agreement with independently verified ground-truth labels.
@@ 5.1 Quantitative Performance
Table 2: Performance metrics for Gandahar Market (median across folds).
Imagery Method AUPRC Precision Recall F1
4-band Cosine Distance 0.56 0.61 0.53 0.57
CVA 0.58 0.61 0.52 0.56
IR-MAD 0.65 0.66 0.59 0.62
LRC 0.74 0.68 0.75 0.71
8-band Cosine Distance 0.60 0.64 0.55 0.59
CVA 0.62 0.64 0.56 0.60
IR-MAD 0.70 0.71 0.63 0.67
LRC 0.80 0.72 0.79 0.75
Time-series Cosine Distance 0.58 0.62 0.54 0.57
CVA 0.59 0.63 0.55 0.58
IR-MAD 0.66 0.70 0.62 0.66
LRC 0.81 0.74 0.80 0.77
Table 3: Performance metrics for El Fasher.
Imagery Method AUPRC Precision Recall F1
4-band Cosine Distance 0.43 0.51 0.45 0.48
CVA 0.48 0.51 0.44 0.47
IR-MAD 0.50 0.56 0.50 0.53
LRC 0.68 0.62 0.66 0.64
8-band Cosine Distance 0.44 0.48 0.42 0.45
CVA 0.46 0.49 0.43 0.46
IR-MAD 0.49 0.54 0.48 0.51
LRC 0.62 0.57 0.63 0.60
Time-series Cosine Distance 0.40 0.48 0.42 0.45
CVA 0.41 0.49 0.43 0.46
IR-MAD 0.48 0.55 0.50 0.52
LRC 0.65 0.60 0.63 0.61
Tables 26 summarize bootstrapped metrics for each location and imagery configuration. The proposed VAE-based approach, referred to as latent representation change (LRC), consistently outperformed cosine distance, CVA, and IR-MAD across all sites and imagery configurations. Relative improvements in AUPRC over IR-MAD, computed as $\frac{\mathrm{AUPRC}_{\text{LRC}}-\mathrm{AUPRC}_{\text{IR-MAD}}}{\mathrm{AUPRC}_{\text{IR-MAD}}}$, ranged from 14% to 36%, with sites having clearer fire signatures (e.g., Muqrin and Sarafaya) achieving AUPRC values exceeding 0.80. Although these case studies correspond to confirmed conflict events, fire-affected tiles comprise less than 10% of the scene area. The remaining tiles establish the nominal background distribution, representing unburned urban surfaces, vegetation, and other background variability.
Table 4: Performance metrics for Muqrin.
Imagery Method AUPRC Precision Recall F1
4-band Cosine Distance 0.62 0.66 0.59 0.62
CVA 0.63 0.67 0.58 0.62
IR-MAD 0.65 0.64 0.70 0.67
LRC 0.77 0.79 0.82 0.80
8-band Cosine Distance 0.65 0.68 0.62 0.66
CVA 0.67 0.69 0.61 0.65
IR-MAD 0.69 0.67 0.72 0.69
LRC 0.79 0.81 0.83 0.82
Time-series Cosine Distance 0.64 0.69 0.62 0.65
CVA 0.65 0.70 0.63 0.66
IR-MAD 0.71 0.69 0.74 0.71
LRC 0.81 0.83 0.85 0.83
Table 5: Performance metrics for Jaranga.
Imagery Method AUPRC Precision Recall F1
4-band Cosine Distance 0.41 0.49 0.42 0.45
CVA 0.46 0.49 0.41 0.45
IR-MAD 0.52 0.54 0.48 0.51
LRC 0.65 0.60 0.64 0.62
8-band Cosine Distance 0.49 0.55 0.48 0.51
CVA 0.54 0.55 0.48 0.51
IR-MAD 0.63 0.63 0.57 0.60
LRC 0.72 0.67 0.69 0.68
Time-series Cosine Distance 0.46 0.52 0.45 0.48
CVA 0.47 0.53 0.46 0.49
IR-MAD 0.55 0.60 0.54 0.57
LRC 0.74 0.68 0.71 0.69
Across all sites, metrics, and imagery types, the 95% confidence intervals were generally narrow, with a median span of 0.01 to 0.04, reflecting low to moderate variability in the bootstrapped estimates for most scenes, although resampling within scenes may underestimate uncertainty if observations are not fully independent. In the highly heterogeneous urban setting of El Fasher, wider intervals were observed, consistent with its more complex scene structure.
Table 6: Performance metrics for Sarafaya.
Imagery Method AUPRC Precision Recall F1
4-band Cosine Distance 0.64 0.69 0.62 0.65
CVA 0.67 0.70 0.61 0.66
IR-MAD 0.66 0.72 0.66 0.69
LRC 0.78 0.80 0.83 0.81
8-band Cosine Distance 0.64 0.68 0.63 0.65
CVA 0.69 0.71 0.63 0.67
IR-MAD 0.70 0.75 0.69 0.72
LRC 0.80 0.81 0.84 0.83
Time-series Cosine Distance 0.65 0.71 0.64 0.67
CVA 0.66 0.72 0.65 0.68
IR-MAD 0.70 0.77 0.72 0.74
LRC 0.82 0.84 0.86 0.84
Table 7: Cohen’s $d$ effect sizes for LRC against IR-MAD across all metrics.
Place Imagery Cohen’s $d$ Cohen’s $d$ Cohen’s $d$ Cohen’s $d$
(AUPRC) (Precision) (Recall) (F1)
Gandahar Market 4-band 1.15 0.45 1.30 1.10
8-band 1.18 0.48 1.32 1.12
Time-series 1.16 0.46 1.31 1.11
El Fasher 4-band 1.10 0.48 1.25 1.05
8-band 0.75 0.35 0.85 0.70
Time-series 1.05 0.45 1.15 0.98
Muqrin 4-band 1.25 0.50 1.35 1.20
8-band 1.27 0.52 1.36 1.22
Time-series 1.26 0.51 1.35 1.21
Jaranga 4-band 1.05 0.42 1.20 1.00
8-band 1.08 0.45 1.22 1.02
Time-series 1.06 0.44 1.21 1.01
Sarafaya 4-band 1.20 0.48 1.30 1.15
8-band 1.22 0.50 1.31 1.16
Time-series 1.21 0.49 1.30 1.15
In addition to absolute performance metrics, in Table 7 we quantified the magnitude of improvement of the VAE-based approach over IR-MAD using Cohen’s $d$, since other models generally exhibited lower performance. Following the paired tile-level experimental design, effect sizes were computed using the standard deviation of paired differences across corresponding $32\times 32$ tiles within each bootstrap sample. Effect sizes ranged from 0.35 to 1.36 across metrics, with the largest values observed for Recall ($d\approx 1.30$), reflecting the framework’s design to prioritize detection of fire-affected areas while maintaining operationally viable precision. Paired Wilcoxon signed-rank tests indicated that these differences were statistically significant ($p<0.01$), consistent with the magnitude of the estimated effect sizes.